Two identical supersonic flights are approaching each other at the same speed. Each flight observes the frequency of the other flight as three times as much as its own frequency. If the speed of sound in air is 330 m/s. calculate the speed of each flight?
Answer #1

V-V f=fo (v tvo s as vous 3 for fo ( 330 tu u 330 - V. ( 330 - V.) 3 330 + Vo 990 -3 Vos 330 + vo 660- a vo vo= 165 m/s. A